OpenStudy (jenniferjuice):

SERIOUSLY GUYS AGINDAJSBHNF HALPPPP Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. y = a(x-h)² + k a= 4 -(-4) h=8 k=16 y=4(x-8)²+16 Using complete sentences, explain how to convert your birthday function into standard form. 8/16/99

OpenStudy (anonymous):

i cans see the graph or whatever it was but i love your profile pic its kellin :D

OpenStudy (anonymous):

seriously.....do your own work.

OpenStudy (jenniferjuice):

@cookies8701 :D thanks kellin rules c:

OpenStudy (jenniferjuice):

@cookies8701 there is no graph its just that question and im confused >.<

OpenStudy (anonymous):

yas i ship kellic but i have no idea how to do it is there anything else i can help you with sorry

OpenStudy (jenniferjuice):

:( okay D:

OpenStudy (anonymous):

is there anything else you need help on

OpenStudy (anonymous):

im so confused...

OpenStudy (anonymous):

with the question?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ppl are talking about graphs and crap and their isn't any graphs

OpenStudy (anonymous):

ya i thought there was a graph but there isnt one

OpenStudy (anonymous):

oh... k

OpenStudy (jenniferjuice):

its a function i have to put it in standard form but idk what a b and c are supposed to be.

OpenStudy (anonymous):

ooo i dunno...

OpenStudy (anonymous):

a b and c are most likely the month the day and year

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

:) yay thanks for the medal

OpenStudy (jenniferjuice):

do you remember hat standard form is ._.

OpenStudy (anonymous):

does that help

OpenStudy (jenniferjuice):

not really

OpenStudy (jenniferjuice):

@ganeshie8

OpenStudy (jenniferjuice):

reread the question im pretty sure its good now. @cookies8701 @ganeshie8

ganeshie8 (ganeshie8):

8/16/99 \[\large y = a(x-h)^2+k\] you may choose \(a=2\) \(h\) = your DOB month = 8 \(k\) = your DOB day = 16 so the equation becomes : \[\large y = 2(x-8)^2+16\]

ganeshie8 (ganeshie8):

you need to change it to standard form : \(y =ax^2+bx+c\)

ganeshie8 (ganeshie8):

familiar with (a-b)^2 formula ?

OpenStudy (jenniferjuice):

i think

ganeshie8 (ganeshie8):

apply that formula and simplify the previous equation

ganeshie8 (ganeshie8):

can you ?

OpenStudy (jenniferjuice):

wait whats b and c?

OpenStudy (jenniferjuice):

@ganeshie8

OpenStudy (jenniferjuice):

dsjgvafjbhdas